In concentration of aqueous solutions find mass fraction and volume fraction using mole fraction 0.7304, specific gravity 0.6268, and molecular weight solute 42.00.
1
Expert's answer
2021-10-07T03:44:25-0400

Volum fraction = 0.7304

Specific gravity of solute=0.6268

Molecular weight of solute=42.00

"Volume fraction=\frac{Mass \; fraction}{Specific \; gravity}\\

Mass fraction= Volum fraction \times Specific gravity \\

= 0.7304 \times 0.6268 \\

Mass fraction= 0.46 \\

Mole fraction = \frac{Mass \; fraction}{Mplecular weight}\\

= \frac{0.46}{42}"


Mole fraction= 0.01095